OpenStudy (anonymous):

I need help proving one more identity: (cos(x) / (1 + cos(x))) + cos(x) / (1 - cos(x)) = 2cot(x) csc(x) Will give medal & fan. I have all the step-by-step, but I need the reasons. I'll post the steps in the comments.

OpenStudy (hba):

I think you should use that sheet now.

OpenStudy (anonymous):

This one is still confusing me though /: Can I tell you what I think the reason is and you tell me if it's correct or incorrect?

OpenStudy (hba):

Try it once and then get your answers confirmed :)

OpenStudy (hba):

Oh yeah sure :)

hero (hero):

What sheet?

hero (hero):

lol, really?

OpenStudy (hba):

or the other one which i game her.She doesn't know the formulas .-.

OpenStudy (hba):

gave*

OpenStudy (anonymous):

Take cos x outside and the you have the eqn... Cosx[(1/1+cosx)+(1/1-cosx)] Solve this and you get Cosx[2/(1-cosx^2)] Then you have Cosx[2/sinx^2] By simplifying....you get the solution

hero (hero):

First figure out how to combine the fractions together. Find the least common demominator.

OpenStudy (anonymous):

The first step is: (cos(x) / (1 + cos(x))) + (cos(x) / (1 - cos(x))) = ((cos(x)(1 - cos(x)) + cos(x)(1 + cos(x))) / (1 + cos(x))(1 - cos(x)) I just can't figure out what would make it all cos(x) /: I'm finding csc = 1/sin and cot = 1/tan

hero (hero):

You mean, you have all the steps already but you still can't figure it out?

OpenStudy (anonymous):

I need to put reasons with the steps...

OpenStudy (anonymous):

Yeah I'm kind of stupid, you got me. And I'm not being sarcastic.

hero (hero):

What kind of reasons are they expecting you to give for this?

OpenStudy (hba):

lulz

OpenStudy (anonymous):

I guess the basic identities they're using to get those steps? I'm not really sure

OpenStudy (anonymous):

You have to understand I don't get a lot of instruction with this. My teacher does nothing for me except grade my work. I teach myself everything, without guidance. And this is quite difficult for me.

hero (hero):

I think for the first step, since you are multiplying both fractions by the equivalent of 1 you can say \(\dfrac{\cos(x) + 1}{\cos(x) + 1} = 1\) and \(\dfrac{1 - \cos(x) }{1 - \cos(x) } = 1\)

OpenStudy (anonymous):

Are you sure? I mean I know you know better than I do lol but all the cos's are confusing me

hero (hero):

Multiplying each of the fractions by the equivalent of one is the only way to get the LCD necessary to combine the fractions.

OpenStudy (anonymous):

This is the problem I'm working on w/ the steps included.

hero (hero):

Yep, it's exactly what I said.

OpenStudy (anonymous):

So what would I do for the second step?

OpenStudy (anonymous):

The second step would be to factor?

OpenStudy (nincompoop):

no. I misread

OpenStudy (anonymous):

Ohh okay

OpenStudy (nincompoop):

simplify. cos^2 x gets cancel out you have + and -

OpenStudy (nincompoop):

the denominator uses pythagorean

hero (hero):

In the second step, cos(x) is distributed

hero (hero):

since \(\cos(x)(1 + \cos(x)) = \cos(x) + \cos^2x\)

hero (hero):

and \(\cos(x)(1 - \cos(x)) = \cos(x) - \cos^2x\)

hero (hero):

@nincompoop, she already has the steps.

hero (hero):

She needs the "reasons" for each step.

OpenStudy (anonymous):

I am so confused... I got step 2 from you but now I'm stuck on step three lol

hero (hero):

I'm surprised you can't figure out how to get to step 3 from step 2

hero (hero):

What's \(\cos^2(x) - \cos^2(x)\)?

OpenStudy (anonymous):

0

hero (hero):

What's \(\cos(x) + \cos(x)\)?

OpenStudy (anonymous):

2cos(x)

hero (hero):

Okay so there you go.

hero (hero):

Oh and also for the denominator

OpenStudy (anonymous):

And I guess 1 - cos(x) is sin^2 (x)?

hero (hero):

Pythagorean Identity rule applies

hero (hero):

Listing reasons is almost kind of pointless here.

OpenStudy (anonymous):

I know but if I don't do it I get a 0 on it so x_x

OpenStudy (anonymous):

For the fourth step is sin^2 (x) divided into the numerator to become 2(cos(x) / sin(x))(1/sin(x)) ?

hero (hero):

The fourth step is more of what happens when you write \(\dfrac{4}{a^2}\) as \(\left(\dfrac{4}{a}\right)\left(\dfrac{1}{a}\right)\)

OpenStudy (anonymous):

... I don't understand that /:

OpenStudy (anonymous):

oh wait.. yeah I do but what would be a?

OpenStudy (anonymous):

Oh wait I think I got it, thanks